# Ehrhart quasi-polynomials via Barnes polynomials and discrete moments of parallelepipeds

### v1

@@ 1 Introduction, and statements of the main results
valid for almost all $z\in\mathbb{C}^{2}$ (see Example 5). Such identities appear to be new. The integer $12$ is easy to explain - it arises naturally from the second classical Bernoulli number $B_{2}=\frac{1}{6}$. Similarly, we have a corresponding identity for all $3$-dimensional unimodular polytopes:
$24=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{1}{2}\right).$
@@
There is now a vast literature on Ehrhart polynomials and quasi-polynomials, and the reader may consult, for example, the books [4] [9] [21] [25] [37] [38], and papers [2] [5] [10] [12] [14] [15] [16] [17] [18] [19] [20] [22] [23] [24] [26] [29] [30] [31] [33] [34] [35] [39]. We remark that given a triangulation of any rational polytope into partially-open simplices, one can obtain similar formulas as in this paper. However, for general rational polytopes, finding an explicit triangulation is computationally expensive, even for general $3$-dimensional polytopes: some of its vertices might have $2^{100}$ edges, for example.
2 Known properties of Barnes polynomials
where the $B_{m}(x)$ are the classical Bernoulli polynonomials, ${\mathbf{a}}:=(a_{1},\dots,a_{d})\in\mathbb{C}^{d}$, and the indices are nonnegative integers $m_{j}\geq 0$ (see Bayad and Kim [7, Theorem 2]). We see that if all of the $a_{j}$’s are rational numbers, then $B_{k}(t,{\mathbf{a}})$ is a polynomial in $t$, with rational coefficients. From (9) and (10), it follows at once that $(a_{1}a_{2}\cdots a_{d})B_{k}(t,{\mathbf{a}})\in\mathbb{Q}[a_{1},\dots,a_{d}]$. Also, it’s clear that the coefficients of $B_{k}(t,{\mathbf{a}})$, as a polynomial in $t$, are symmetric functions of $a_{1},\dots,a_{d}$. The Barnes polynomials enjoy the following properties.
[reciprocity law]
$(-1)^{k}B_{k}(-t,{\mathbf{a}})=B_{k}(t+a_{1}+\cdots+a_{d},{\mathbf{a}}),$ (34)
where ${\mathbf{a}}:=(a_{1},\dots,a_{d})^{T}$.
[homogeneity]
$B_{k}(\lambda t,\lambda{\mathbf{a}})=\lambda^{k-d}B_{k}(t,{\mathbf{a}}),$ (35)
for each $k\geq 1$, and for any fixed nonzero $\lambda\in\mathbb{C}$. Here by definition $\lambda{\mathbf{a}}:=(\lambda a_{1},\dots,\lambda a_{d})$. See [7] for proofs and more background on Barnes polynomials.
Using identity (9), we list the first few Barnes numbers with ${\mathbf{a}}\in\mathbb{C}^{d}$:
$B_{0}({\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}}.$ (36)
$(a_{1}\cdots a_{d})B_{2}({\mathbf{a}})=\frac{1}{6}\sum_{k=1}^{d}a_{k}^{2}+\frac{1}{2}\sum_{1\leq i<j\leq d}a_{i}a_{j}.$ (38)
$(a_{1}\cdots a_{d})B_{3}({\mathbf{a}})=-\frac{1}{4}\sum_{1\leq i\not=j\leq d}a_{i}^{2}a_{j}-\frac{1}{8}\sum_{1\leq i<j<k\leq d}a_{i}a_{j}a_{k}.$ (39)
In addition, using the identities (10) and (9), we list the first few Barnes polynomials with ${\mathbf{a}}\in\mathbb{C}^{d}$:
$B_{0}(t,{\mathbf{a}})=\frac{1}{a_{1}\cdots a_{d}},$ (40)
$B_{1}(t,{\mathbf{a}})$ $=\binom{1}{0}B_{0}({\mathbf{a}})t+\binom{1}{1}B_{1}({\mathbf{a}})$ (41)
$=\frac{1}{a_{1}\cdots a_{d}}t-\frac{1}{2}\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}.$ (42)
$B_{2}(t,{\mathbf{a}})=\binom{2}{0}B_{0}({\mathbf{a}})t^{2}+\binom{2}{1}B_{1}({\mathbf{a}})t+\binom{2}{2}B_{2}({\mathbf{a}})$ (43)
$=\frac{1}{a_{1}\cdots a_{d}}t^{2}-\Big(\frac{a_{1}+a_{2}+\cdots+a_{d}}{a_{1}\cdots a_{d}}\Big)t+\Big(\frac{1}{6}\sum_{k=1}^{d}\frac{a_{k}^{2}}{a_{1}\cdots a_{d}}+\frac{1}{2}\sum_{1\leq i<j\leq d}\frac{a_{i}a_{j}}{a_{1}\cdots a_{d}}\Big).$ (44)
Finally, for the purposes of explicitly computing some Ehrhart polynomials, we list the first few relevant Barnes polynomials, for each dimension $d=1,2,3$, derived using the relations (9), (10), (36), (37), (38), and (39):
$B_{1}(t,(a_{1}))=\frac{1}{a_{1}}t-\frac{1}{2}.$ (45)
$B_{2}\left(t,\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right)\right)=t^{2}\left(\frac{1}{a_{1}a_{2}}\right)-t\left(\frac{1}{a_{2}}+\frac{1}{a_{1}}\right)+\frac{1}{6}\left(\frac{a_{2}}{a_{1}}+\frac{a_{1}}{a_{2}}\right)+\frac{1}{2}.$ (46)
$B_{3}\left(t,\left(\begin{smallmatrix}a_{1}\\ \\ a_{2}\\ \\ a_{3}\end{smallmatrix}\right)\right)=\binom{3}{0}B_{0}({\mathbf{a}})t^{3}+\binom{3}{1}B_{1}({\mathbf{a}})t^{2}+\binom{3}{2}B_{2}({\mathbf{a}})t+\binom{3}{3}B_{3}({\mathbf{a}})$ (47)
$=\frac{1}{a_{1}a_{2}a_{3}}t^{3}-\frac{3}{2}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}}\right)t^{2}$ (48)
$+\frac{3}{a_{1}a_{2}a_{3}}\left(\frac{1}{6}(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})+\frac{1}{2}(a_{1}a_{2}+a_{1}a_{3}+a_{2}a_{3})\right)t$ (49)
$+\frac{1}{a_{1}a_{2}a_{3}}\Big(-\frac{1}{4}\sum_{1\leq i\not=j\leq 3}a_{i}^{2}a_{j}-\frac{1}{8}\sum_{1\leq i<j<k\leq 3}a_{i}a_{j}a_{k}\Big)$ (50)
$=\frac{1}{a_{1}a_{2}a_{3}}t^{3}-\frac{3}{2}\left(\frac{a_{1}+a_{2}+a_{3}}{a_{1}a_{2}a_{3}}\right)t^{2}$ (51)
$+\frac{1}{2}\left(\left(\frac{a_{1}}{a_{2}a_{3}}+\frac{a_{2}}{a_{1}a_{3}}+\frac{a_{3}}{a_{1}a_{2}}\right)+3\left(\frac{1}{a_{1}}+\frac{1}{a_{2}}+\frac{1}{a_{3}}\right)\right)t$ (52)
$-\frac{1}{4}\left(\frac{a_{1}}{a_{2}}+\frac{a_{1}}{a_{3}}+\frac{a_{2}}{a_{1}}+\frac{a_{2}}{a_{3}}+\frac{a_{3}}{a_{1}}+\frac{a_{3}}{a_{2}}\right)-\frac{1}{8}.$ (53)
3 Dimension 1: intervals
@@
and here the vector of linear forms is just ${\mathbf{a}}_{v}=z$ for the vertex $v:=a$, and ${\mathbf{a}}_{v}=-z$ for the vertex $v:=b$. In dimension $1$ all polytopes are unimodular, so by Corollary 2 we have:
@@
$=-\sum_{v\in\rm{V}}B_{1}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)-\sum_{v\in\rm{V}}B_{0}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\mathbb{Z}^{d}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}$
@@
4 Dimension 2: polygons
Suppose we are given any integer polygon $\mathcal{P}$}, defined as the convex hull of its vertices $\{v_{1},\cdots v_{N}\}$. At each vertex $v$, we are given its edge vectors $w_{1}(v),w_{2}(v)$. In the ensuing formula, we define $c_{v}:=\frac{1}{2}(w_{1}(v)+w_{2}(v))$, and the indicator function $1_{\mathbb{Z}^{2}}(x):=\begin{cases}1&\text{if $x\in\mathbb{Z}^{2}$}\\ 0&\text{otherwise}.\end{cases}$
An explicit description for the Ehrhart polynomial of $\mathcal{P}$}, using Corollary 1, is as follows:
$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+\left(\frac{1}{2}\sum_{v\in V}\langle v,z\rangle\left(\frac{1}{\langle w_{1}(v),z\rangle}+\frac{1}{\langle w_{2}(v),z\rangle}\right)\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)\right)t+1,$ (55)
valid for almost all $z\in\mathbb{C}^{2}$ and $t\in\mathbb{Z}_{>0}$. To show how this follows from Theorem 1, we first recall the definition ${\mathbf{a}}_{v}:=\left(\begin{smallmatrix}\langle w_{1}(v),z\rangle\\ \langle w_{2}(v),z\rangle\end{smallmatrix}\right)$ at each vertex $v$ of $\mathcal{P}$}. From Theorem 1, we have:
@@ 4 Dimension 2: polygons
To summarize, we have:
@@
valid for almost all $z\in\mathbb{C}^{2}$ and $t\in\mathbb{Z}_{>0}$. We used Corollary 5 for the constant term in (56), and equation (111) for the area of the polygon. Recalling that $c_{v}:=\frac{1}{2}(w_{1}(v)+w_{2}(v))$, the center of mass of the closure of the parallelogram $\Pi_{v}$, it’s an easy exercise that summing all the points in a fundamental parallelepiped gives us:
$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}p=\big(\operatorname{vol}\Pi_{v}+1-1_{\mathbb{Z}^{2}}(c_{v})\big)\ c_{v}.$ (57)
(hint: translate $\Pi_{v}\cap\mathbb{Z}^{2}$ to the origin by moving its center of mass, and notice that the new non-zero points cancel in paris). So we have:
$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}\langle p,z\rangle-\operatorname{vol}\Pi_{v}\langle c_{v},z\rangle=(1-1_{\mathbb{Z}^{2}}(c_{v}))\langle c_{v},z\rangle.$
@@
$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+t\sum_{v\in V}\frac{\langle v,z\rangle\langle c_{v},z\rangle}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)+1$
$={\rm area}({\mathcal{P}})t^{2}+\left(\frac{1}{2}\sum_{v\in V}\langle v,z\rangle\left(\frac{1}{\langle w_{1}(v),z\rangle}+\frac{1}{\langle w_{2}(v),z\rangle}\right)\left(1-1_{\mathbb{Z}^{2}}(c_{v})\right)\right)t+1,$
as claimed.
@@
We may also observe that for any lattice flow, the orbit of any integer point in $\Pi_{v}$ produces a closed geodesic that has the same length as any other integer point. The reason for these equal length geodesics is easy, and arises from the finite abelian group $\mathbb{Z}^{d}/\Lambda$ structure, where $\Lambda$ is the integer sublattice generated by the vertices of $\Pi_{v}$. Thus, for example in Figure 2, the total length of the red geodesics equals the total length of the blue geodesics, because the blue geodesics arise from an index $2$ subgroup of $\mathbb{Z}^{2}/\Lambda$ here.
$\square$
@@
In $\mathbb{R}^{2}$, just for fun let’s retrive the Ehrhart polynomial for the standard triangle ${\mathcal{P}}:=\operatorname*{conv}\{\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)\}$, using Barnes polynomials. We’ll use Corollary 2 of Theorem 1, since the standard triangle is a unimodular polygon. Recalling that the only Barnes polynomials that occur in dimension $2$ are:
@@
Hence by definition, we have ${\mathbf{a}}_{v_{1}}:=\Big(\langle z,\omega_{1}(v_{1})\rangle,\langle z,\omega_{2}(v_{1})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}\\ z_{2}\end{smallmatrix}\right)$. Similarly, by definition we have ${\mathbf{a}}_{v_{2}}:=\Big(\langle z,\omega_{1}(v_{2})\rangle,\langle z,\omega_{2}(v_{2})\rangle\Big)^{T}=\left(\begin{smallmatrix}-z_{1}+z_{2}\\ -z_{1}\end{smallmatrix}\right)$, and
${\mathbf{a}}_{v_{3}}:=\Big(\langle z,\omega_{1}(v_{3})\rangle,\langle z,\omega_{2}(v_{3})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}-z_{2}\\ -z_{2}\end{smallmatrix}\right)$. By Corollary 2 , we have
@@
5 The unit cube in dimension d
We consider the $d$-dimensional unit cube $\square:=[-1,1]^{d}$, another unimodular polytope. Although it is trivial to compute the Ehrhart polynomial in this case, namely $L_{\square}(t)=(2t+1)^{d}$, the verification will shed some light on our use of the Barnes polynomials, and even gives new identities (equation (61) below) for the discrete hypercube $\{-1,1\}^{d}$.

### v2

@@ 1.4 Smooth polytopes
valid for almost all $z\in\mathbb{C}^{2}$ (see Example 5). The integer $12$ is easy to explain - it arises naturally from the second classical Bernoulli number $B_{2}=\frac{1}{6}$. Similarly, we easily get a corresponding identity for all $3$-dimensional smooth polytopes:
$24=\sum_{v\in\rm{V}}\left(\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{2}(v),z\rangle}+\tfrac{\langle w_{1}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{2}(v),z\rangle}{\langle w_{3}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{1}(v),z\rangle}+\tfrac{\langle w_{3}(v),z\rangle}{\langle w_{2}(v),z\rangle}+3\right),$
@@
valid for almost all $z\in\mathbb{C}^{3}$ (see Example 6). Such identities appear to be new. There is now a vast literature on Ehrhart polynomials and quasi-polynomials, and the reader may consult, for example, the books [5] [10] [25] [29] [42] [43], and papers [2] [3] [6] [11] [13] [15] [16] [17] [18] [19] [20] [21] [22] [23] [24] [26] [27] [28] [30] [33] [34] [35] [37] [38] [39] [40] [44]. We remark that given a triangulation of any rational polytope into partially-open simplices, one can obtain similar formulas to those we obtained in Theorem 1 (and consequently the other theorems here), for its Ehrhart quasi-polynomial and its discrete moments. However, for general rational polytopes, finding an explicit triangulation is computationally expensive, even for general $3$-dimensional polytopes: some of its vertices might have $2^{100}$ edges, for example.
2 Small dimensions
Here we show how to specialize the formula of Theorem 1 to obtain the classical results in dimensions $1$ and $2$.
2.1 Dimension 1: intervals
@@
and here the vector of linear forms is just ${\mathbf{a}}_{v}=z$ for the vertex $v:=a$, and ${\mathbf{a}}_{v}=-z$ for the vertex $v:=b$. In dimension $1$ all polytopes are smooth, so by Corollary 5 we have:
@@
$=-\sum_{v\in\rm{V}}B_{1}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)-\sum_{v\in\rm{V}}B_{0}\Big(t\langle v,z\rangle,{\mathbf{a}}_{v}\Big)\sum_{p\in\mathbb{Z}-tv\;(\hskip-2.84544pt\bmod\Pi_{v})}{\langle p,z\rangle}$
@@
2.2 Dimension 2: polygons
Suppose we are given any integer polygon $\mathcal{P}$}, defined as the convex hull of its vertices $\{v_{1},\cdots v_{N}\}$. At each vertex $v$, we are given its primitive edge vectors $w_{1}(v),w_{2}(v)$. An explicit description for the Ehrhart polynomial of $\mathcal{P}$}, using Corollary 1, is as follows:
$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}-\frac{1}{2}\sum_{v\in V}\left(\frac{\langle v,z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle v,z\rangle}{\langle w_{2}(v),z\rangle}\right)t+1,$ (36)
valid for almost all $z\in\mathbb{C}^{2}$ and $t\in\mathbb{Z}_{>0}$. To show how this follows from Corollary 1, we first recall the definition ${\mathbf{a}}_{v}:=\left(\begin{smallmatrix}\langle w_{1}(v),z\rangle\\ \langle w_{2}(v),z\rangle\end{smallmatrix}\right)$ at each vertex $v$ of $\mathcal{P}$}. By Corollary 1, we have:
@@ 2.2 Dimension 2: polygons
To simplify notation, we define
$c_{v}:=\frac{1}{2}(w_{1}(v)+w_{2}(v)),$
and we summarize:
@@
valid for almost all $z\in\mathbb{C}^{2}$ and $t\in\mathbb{Z}_{>0}$. We used Corollary 2 for the constant term in (37), and equation (99) for the area of the polygon. It’s an easy exercise that summing all the points in a $2$-dimensional fundamental parallelepiped gives us:
$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}p=\big(\operatorname{vol}\Pi_{v}-1\big)\ c_{v}.$ (38)
[hint: translate $\Pi_{v}\cap\mathbb{Z}^{2}$ to the origin by (its center of mass) $c_{v}$, and notice that the new non-zero points cancel in pairs]. So we have:
$\sum_{p\in\Pi_{v}\cap\mathbb{Z}^{2}}\langle p,z\rangle-\operatorname{vol}\Pi_{v}\langle c_{v},z\rangle=-\langle c_{v},z\rangle.$
@@
$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}-t\sum_{v\in V}\frac{\langle v,z\rangle\langle c_{v},z\rangle}{\langle w_{1}(v),z\rangle\langle w_{2}(v),z\rangle}+1$ (39)
$={\rm area}({\mathcal{P}})t^{2}-\frac{1}{2}\sum_{v\in V}\left(\frac{\langle v,z\rangle}{\langle w_{1}(v),z\rangle}+\frac{\langle v,z\rangle}{\langle w_{2}(v),z\rangle}\right)t+1,$ (40)
using the definition $\langle c_{v},z\rangle=\frac{1}{2}\left(\langle w_{1}(v),z\rangle+\langle w_{2}(v),z\rangle\right)$.
Finally, let’s see how to translate the coordinate-free formulation (40) into the standard known result on Ehrhart polynomials. For any $x\in\mathbb{R}^{d}$, we define $\gcd(x):=|\gcd(x_{1},\dots,x_{d})|$. We consider any four consecutive vertices of $\mathcal{P}$}, say $v_{1},v_{2},v_{3},v_{4}$. To give the primitive edge vectors for $v_{2}$, we define $g_{21}:=\gcd(v_{1}-v_{2})$, which is the number of integer points on the half-open edge joining $v_{1}$ and $v_{2}$. We note also that, for example, $g_{23}=g_{32}$. Then $w_{1}(v_{2})=\frac{1}{g_{21}}(v_{1}-v_{2})$, and similarly $w_{2}(v_{2})=\frac{1}{g_{23}}(v_{3}-v_{2})$. In the summation of (40), we focus on just the two vertices $v_{2}$ and $v_{3}$, and we consider their contribution to the sum:
$\left(\frac{\langle v_{2},z\rangle}{\langle w_{1}(v_{2}),z\rangle}+\frac{\langle v_{2},z\rangle}{\langle w_{2}(v_{2}),z\rangle}\right)+\left(\frac{\langle v_{3},z\rangle}{\langle w_{1}(v_{3}),z\rangle}+\frac{\langle v_{3},z\rangle}{\langle w_{2}(v_{3}),z\rangle}\right)$ (41)
$=\frac{g_{21}\langle v_{2},z\rangle}{\langle v_{1}-v_{2},z\rangle}+\left(\frac{g_{23}\langle v_{2},z\rangle}{\langle v_{3}-v_{2},z\rangle}+\frac{g_{32}\langle v_{3},z\rangle}{\langle v_{2}-v_{3},z\rangle}\right)+\frac{g_{34}\langle v_{3},z\rangle}{\langle v_{4}-v_{3},z\rangle}$ (42)
$=\frac{g_{21}\langle v_{2},z\rangle}{\langle v_{1}-v_{2},z\rangle}-g_{23}+\frac{g_{34}\langle v_{3},z\rangle}{\langle v_{4}-v_{3},z\rangle}.$ (43)
Once we consider the full summation in (40), we therefore see that the simplification in pairs gives us the classical result:
$L_{\mathcal{P}}(t)$ $={\rm area}({\mathcal{P}})t^{2}+\frac{1}{2}t\sum_{j=1}^{N}\gcd\left(v_{j}-v_{j+1}\right)+1,$ (44)
where $N$ is the number of vertices of $\mathcal{P}$}, and by definition $v_{N+1}:=v_{1}$.
@@
We may also observe that for any lattice flow, the orbit of any integer point in $\Pi_{v}$ produces a closed geodesic that has the same length as any other integer point. The reason for these equal length geodesics is easy, and arises from the finite abelian group $\mathbb{Z}^{d}/\Lambda$ structure, where $\Lambda$ is the integer sublattice generated by the vertices of $\Pi_{v}$. Thus, for example in Figure 2, there are exactly two closed geodesics, and the total length of the red geodesic equals the total length of the blue geodesic, because here the blue geodesic arises from an index $2$ subgroup of $\mathbb{Z}^{2}/\Lambda$. $\square$
@@
In $\mathbb{R}^{2}$, just for fun let’s retrieve the Ehrhart polynomial for the standard triangle ${\mathcal{P}}:=\operatorname*{conv}\{\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right),\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)\}$, using Barnes polynomials. We’ll use Corollary 5 of Theorem 1, since the standard triangle is a smooth polygon. We recall that the only Barnes polynomials that occur in dimension $2$ are:
@@
Therefore by definition, we have ${\mathbf{a}}_{v_{1}}:=\Big(\langle z,\omega_{1}(v_{1})\rangle,\langle z,\omega_{2}(v_{1})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}\\ z_{2}\end{smallmatrix}\right)$. Similarly ${\mathbf{a}}_{v_{2}}:=\Big(\langle z,\omega_{1}(v_{2})\rangle,\langle z,\omega_{2}(v_{2})\rangle\Big)^{T}=\left(\begin{smallmatrix}-z_{1}+z_{2}\\ -z_{1}\end{smallmatrix}\right)$, and ${\mathbf{a}}_{v_{3}}:=\Big(\langle z,\omega_{1}(v_{3})\rangle,\langle z,\omega_{2}(v_{3})\rangle\Big)^{T}=\left(\begin{smallmatrix}z_{1}-z_{2}\\ -z_{2}\end{smallmatrix}\right)$. By Corollary 5 , we have
@@
2.3 A d-dimensional cube
We consider the $d$-dimensional cube $\square:=[-1,1]^{d}$, another smooth polytope. Although it is trivial to compute the Ehrhart polynomial in this case, namely $L_{\square}(t)=(2t+1)^{d}$, the verification will shed some light on our use of the Barnes polynomials, and even gives new identities (equation (48) below) for the discrete hypercube $\{-1,1\}^{d}$.